OpenStudy (anonymous):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units?

OpenStudy (anonymous):

please help me guys!

OpenStudy (anonymous):

well think what is the formula to find the perimeter?

OpenStudy (anonymous):

a+b+c?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

P=2(l+w)

OpenStudy (anonymous):

i dont remember

OpenStudy (anonymous):

so now do the math with the formula

OpenStudy (anonymous):

can you give me the answer then we discuss it?

OpenStudy (texaschic101):

perimeter of rectangle is : P = 2L + 2W start subbing in : P = 64 L = 20 W = 11 if it comes out correct, then it can be a solution if it comes out incorrect, then it cannot

OpenStudy (anonymous):

here are my options

OpenStudy (anonymous):

No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64 No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64 Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32

OpenStudy (anonymous):

20+11*2=62

OpenStudy (anonymous):

no option has 62 as an answer

OpenStudy (texaschic101):

ok...lets sub P = 2L + 2W 64 = 2(20) + 2(11) 64 = 40 + 22 64 = 62....it cannot be true

OpenStudy (anonymous):

the= with a line through it means not = to. get it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so whats the final answer from the options?

OpenStudy (anonymous):

I would say A

OpenStudy (anonymous):

ok thx!!!!

OpenStudy (anonymous):

no problem do you need more help?

OpenStudy (texaschic101):

It doesn't make much sense....x and y do not have to equal 64.....2x + 2y have to equal 64

OpenStudy (anonymous):

so whats the answer? texaschic

OpenStudy (anonymous):

yes buttons i could use some more help.

OpenStudy (anonymous):

k what do ya need

OpenStudy (texaschic101):

I think it is B

OpenStudy (triciaal):